# ASiM: Modeling and Analyzing Inference Accuracy of SRAM-Based Analog CiM Circuits

### v3

@@ Abstract
SRAM-based Analog Compute-in-Memory (ACiM) demonstrates promising energy efficiency for deep neural network (DNN) processing. Nevertheless, efforts to optimize efficiency frequently compromise accuracy, and this trade-off remains insufficiently studied due to the difficulty of performing full-system validation. Specifically, existing simulation tools rarely target SRAM-based ACiM and exhibit inconsistent accuracy predictions, highlighting the need for a standardized, SRAM CiM circuit-aware evaluation methodology. This paper presents ASiM, a simulation framework for evaluating inference accuracy in SRAM-based ACiM systems. ASiM captures critical effects in SRAM based analog compute in memory systems, such as ADC quantization, bit parallel encoding, and analog noise, which must be modeled with high fidelity due to their distinct behavior in charge domain architectures compared to other memory technologies. ASiM supports a wide range of modern DNN workloads, including CNN and Transformer-based models such as ViT, and scales to large-scale tasks like ImageNet classification. Our results indicate that bit-parallel encoding can improve energy efficiency with only modest accuracy degradation; however, even 1 LSB of analog noise can significantly impair inference performance, particularly in complex tasks such as ImageNet. To address this, we explore hybrid analog–digital execution and majority voting schemes, both of which enhance robustness without negating energy savings. ASiM bridges the gap between hardware design and inference performance, offering actionable insights for energy-efficient, high-accuracy ACiM deployment. Code is available at https://github.com/Keio-CSG/ASiM
@@ I Introduction
Deep learning has brought revolutionary changes to a wide range of fields, not only in computer vision and natural language processing but also complex tasks spanning different modalities, delivering state-of-the-art performance. Recent advances have shown a scaling law in deep learning models, where performance consistently improves as the size of the model increases [1]. However, executing these large-scale models in traditional von Neumann architectures leads to substantial energy consumption and latency [2], which pose critical challenges, particularly for edge devices with limited resources. Compute-in-Memory (CiM) has emerged as a novel paradigm to overcome these limitations by performing computations directly within the memory arrays  [3]. By reducing data movement between logic and memory, CiM architectures can achieve significant improvements in both energy efficiency and throughput.
@@
The remainder of this paper is organized as follows. Section II reviews related works and provides preliminary concepts on ACiM and inference evaluation metrics. Section III describes the detailed implementation and modeling methodology of the ASiM framework. Section IV presents a comprehensive analysis of how each ACiM design factor impacts DNN inference accuracy and highlights fundamental limitations. Section V provides simple yet effective solutions to improve ACiM inference accuracy. Finally, Section VII concludes the paper.
@@ II-C Evaluation Metrics for Inference Accuracy
To quantify inference degradation from ACiM execution, metrics such as signal-to-quantization noise ratio (SQNR) and compute-signal-to-noise ratio (CSNR) are commonly used [22]. SQNR captures noise from quantization alone, while CSNR accounts for additional analog effects. The original definitions of SQNR and CSNR are given in [23] as:
@@
Here, $y_{n}$ is the ideal full-precision output, and $\hat{y_{n}}$ is the expected output from a real device. Yoshioka [25] extended this approach by introducing analog Gaussian noise into each binary MAC cycle to estimate the actual CSNR of ACiM chips.
While prior studies report high CSNR despite analog noise, our findings indicate these values may not generalize well. Many such estimates assume uniformly distributed activation and weight vectors, leading to inflated signal variance in the numerator of Eq. 3. In actual DNN applications, outputs tend to occupy narrower ranges, reducing true signal power and distorting CSNR interpretation.
@@ III-A Overview
ASiM is implemented within PyTorch to ensure seamless integration with modern DNN workflows. It provides a plug-and-play simulation framework tailored for SRAM-based charge-domain ACiM. For MAC-heavy operations such as CONV, Linear, and Attention, ASiM modules can replace standard PyTorch layers, enabling bit-wise simulations with user-defined circuit parameters.
@@ III-D Noise-Aware Training
We conducted NAT on ResNet-18 and ViT-B-32 using the CIFAR-10 and ImageNet datasets, respectively, with the baseline model accuracy under different training noise intensities value of $\sigma$) shown in Fig. 3. As observed in many previous studies, ResNet models tend to be easier to fine-tune compared to ViT models, and the ImageNet task is generally more challenging than CIFAR-10. Consequently, we selected models optimized for ACiM applications in the following analysis, with a 2.0% accuracy drop for the CIFAR-10 task and around a 3.0% accuracy drop for the ImageNet task.
@@ IV Performance Analysis of ACiM Inference
IV-A Re-evaluating CSNR Under Actual Data Distributions
@@
The self-attention computation demonstrates particularly diverse characteristics:
@@ IV-D Simulation to Silicon Correlation
ADC Error Distribution. To verify that the noise model accurately reflects real ADC error characteristics, we generate random weight and input activation vectors, perform MAC operations on our prototype ACiM chip [30], and measure the resulting voltage using an ADC, comparing it with the ideal output. This process is repeated 30,000 times, and the recorded error distribution is shown in Fig. 12. As depicted, the error exhibits a bell-shaped pattern, where most ADC readings are correct, but occasional errors occur in the range of ±1 to ±3 LSBs. This distribution aligns well with our noise modeling assumption, validating the Gaussian noise model used in ASiM as an accurate representation of the noise encountered in ACiM chips.
@@ VI-C Comparison across State-of-the-Art CiM designs
To evaluate the practical inference performance of SRAM-based ACiM, we compare representative designs that can be accurately modeled within the ASiM framework. All simulations are conducted using a unified 8b/8b ResNet-18 model on the CIFAR-10 benchmark. The resulting inference accuracy, energy efficiency, and area efficiency are normalized to a 22 nm technology node and 0.9 V supply voltage and summarized in Fig. 15. The size of each bubble represents the estimated area efficiency, calculated as the average area per basic CiM cell. Relevant design parameters are extracted from their respective publications and summarized in Table III. For designs where $\text{LSB}_{\text{rms}}$ was not reported, a nominal value of 0.5 LSB was assumed. The designs under evaluation include a range of representative schemes: the standard Bit-Parallel/Bit-Serial (BPBS [24]) architecture for multi-bit DNN processing; the Sparsity-Adaptive (Spar.-Adapt. [13]) scheme, which skips redundant ADC conversions by exploiting activation sparsity; the End-to-End Analog (E2E Analog [40]) pipeline, which maintains analog signal integrity throughout computation; the Differential Switched-Capacitor (Diff.-SC [41]) design, which utilizes differential compute lines to enhance dynamic range and ADC precision; and the Capacitor-Reconfigured CiM (CR-CIM [30]) scheme, which reuses cell capacitors to improve ADC resolution. In addition, CR-CIM is also configured with a reduced 6-bit ADC to demonstrate accuracy degradation, followed by recovery via majority voting. For completeness, the Hybrid CiM (HCiM [42]) and two digital CiM (DCiM) designs [43, 44] are also included as references.

### v4

@@ Abstract
SRAM-based Analog Compute-in-Memory (ACiM) demonstrates promising energy efficiency for Deep Neural Network (DNN) processing. Nevertheless, efforts to optimize efficiency frequently compromise accuracy, and this trade-off remains insufficiently studied due to the difficulty of performing full-system validation. Specifically, existing simulation tools rarely target SRAM-based ACiM and exhibit inconsistent accuracy predictions, highlighting the need for a standardized, SRAM CiM circuit-aware evaluation methodology. This paper presents ASiM, a simulation framework for evaluating inference accuracy in SRAM-based ACiM systems. ASiM captures critical effects in SRAM based analog compute in memory systems, such as ADC quantization, bit parallel encoding, and analog noise, which must be modeled with high fidelity due to their distinct behavior in charge domain architectures compared to other memory technologies. ASiM supports a wide range of modern DNN workloads, including CNN and Transformer-based models such as ViT, and scales to large-scale tasks like ImageNet classification. Our results indicate that bit-parallel encoding can improve energy efficiency with only modest accuracy degradation; however, even 1 LSB of analog noise can significantly impair inference performance, particularly in complex tasks such as ImageNet. To address this, we explore hybrid analog-digital execution and majority voting schemes, both of which enhance robustness without negating energy savings. ASiM bridges the gap between hardware design and inference performance, offering actionable insights for energy-efficient, high-accuracy ACiM deployment. Code is available at https://github.com/Keio-CSG/ASiM
@@ I Introduction
Deep learning has brought revolutionary changes to a wide range of fields, not only in computer vision and natural language processing but also complex tasks spanning different modalities, delivering state-of-the-art performance. Recent advances have shown a scaling law in deep learning models, where performance consistently improves as the size of the model increases [1]. However, executing these large-scale models in traditional von Neumann architectures leads to substantial energy consumption and latency [2], which poses critical challenges, particularly for edge devices with limited resources. Compute-in-Memory (CiM) has emerged as a novel paradigm to overcome these limitations by performing computations directly within the memory arrays  [3]. By reducing data movement between logic and memory, CiM architectures can achieve significant improvements in both energy efficiency and throughput.
@@
The remainder of this paper is organized as follows. Section II reviews related works and provides preliminary concepts on ACiM and inference evaluation metrics. Section III describes the detailed implementation and modeling methodology of the ASiM framework. Section IV presents a comprehensive analysis of how each ACiM design factor impacts DNN inference accuracy and highlights fundamental limitations. Section V provides simple yet effective solutions to improve ACiM inference accuracy. Section VI offers a comprehensive discussion of the ASiM framework alongside state-of-the-art ACiM designs. Finally, Section VII concludes the paper.
@@ II-C Evaluation Metrics for Inference Accuracy
To quantify inference degradation from ACiM execution, metrics such as Signal to Quantization Noise Ratio (SQNR) and Compute Signal to Noise Ratio (CSNR) are commonly used [22]. SQNR captures noise from quantization alone, while CSNR accounts for additional analog effects. The original definitions of SQNR and CSNR are given in [23] as:
@@
Here, $y_{n}$ is the ideal full-precision output, and $\hat{y_{n}}$ is the expected output from a real device. Yoshioka [25] extended this approach by introducing analog Gaussian noise into each binary MAC cycle to estimate the actual CSNR of ACiM chips. While the noise terms in the denominator of Eqs. 2 and 3 can be minimized with careful circuit design, the signal term in the numerator is highly dependent on the specific DNN model and input data. Assuming a uniform distribution across the ADC’s full dynamic range can result in amplified signal variance, leading to an overestimation of CSNR.
@@ III-A Overview
The overview of ASiM framework is shown in Fig. 2. ASiM is implemented within PyTorch to ensure seamless integration with modern DNN workflows. It provides a plug-and-play simulation framework tailored for SRAM-based charge-domain ACiM. For MAC-heavy operations such as CONV, Linear, and Attention, ASiM modules can replace standard PyTorch layers, enabling bit-wise simulations with user-defined circuit parameters.
@@ III-D Noise-Aware Training
We conducted NAT on ResNet-18 and ViT-B-32 using the CIFAR-10 and ImageNet datasets, respectively, with the baseline model accuracy under different training noise intensities value of $\sigma$ shown in Fig. 3. As observed in many previous studies, ResNet models tend to be easier to fine-tune compared to ViT models, and the ImageNet task is generally more challenging than CIFAR-10. Consequently, we selected models optimized for ACiM applications in the following analysis, with a 2.0% accuracy drop for the CIFAR-10 task and around a 3.0% accuracy drop for the ImageNet task.
@@ IV Performance Analysis of ACiM Inference
IV-A Re-consider CSNR with Actual Data Distributions
@@
The attention computation demonstrates particularly diverse characteristics:
@@ IV-D Simulation to Silicon Correlation
ADC Error Distribution. To verify that the noise model accurately reflects real ADC error characteristics, we generate random weight and input activation vectors, perform MAC operations on our prototype ACiM chip [30], and measure the resulting voltage using an ADC, comparing it with the ideal output. This process is repeated 30,000 times, and the recorded error distribution is shown in Fig. 12. As depicted, the error exhibits a bell-shaped pattern, where most ADC readings are correct, but occasional errors occur in the range of 1 to 3 LSBs. This distribution aligns well with our noise modeling assumption, validating the Gaussian noise model used in ASiM as an accurate representation of the noise encountered in ACiM chips.
@@ VI-C Comparison Across State-of-the-Art CiM Designs
To evaluate the practical inference performance of SRAM-based ACiM, we compare representative designs that can be accurately modeled within the ASiM framework. All simulations are conducted using a unified 8b/8b ResNet-18 model on the CIFAR-10 benchmark. The resulting inference accuracy, energy efficiency, and area efficiency are normalized to a 22 nm technology node and 0.9 V supply voltage and summarized in Fig. 15. The size of each bubble represents the estimated area efficiency, calculated as the average area per basic CiM cell. Relevant design parameters are extracted from their respective publications and summarized in Table III. For designs where $\text{LSB}_{\text{rms}}$ was not reported, a nominal value of 0.5 LSB was assumed. The designs under evaluation include a range of representative schemes: the standard Bit-Parallel/Bit-Serial (BPBS [24]) architecture for multi-bit DNN processing; the Sparsity-Adaptive (Spar.-Adapt. [13]) scheme, which skips redundant ADC conversions by exploiting activation sparsity; the End-to-End Analog (E2E Analog [40]) pipeline, which maintains analog signal integrity throughout computation; the Differential Switched-Capacitor (Diff.-SC [41]) design, which utilizes differential compute lines to enhance dynamic range and ADC precision; and the Capacitor-Reconfigured CiM (CR-CIM [30]) scheme, which reuses cell capacitors to improve ADC resolution. In addition, CR-CIM is also configured with a reduced 6-bit ADC to demonstrate accuracy degradation, followed by recovery via majority voting. For completeness, the HCiM (Hybrid [42]) and two DCiM designs [43, 44] are also included as references.